# The stochastic discrete nonlinear Schrödinger equation: microscopic derivation and finite-temperature phase transition

### v1

@@ Abstract
I Introduction
@@
The one-dimensional discrete nonlinear Schrödinger equation (DNSE) is a Hamiltonian system with two conserved quantities that exhibits intriguing features such as localised modes (breathers) and negative temperature states [BaIuLiVu_PR21]. This non-integrable model was initially proposed to describe the dynamics of a group of anharmonic oscillators. Its primary aim was to explain nonlinear localisation phenomena, specifically breathers, i.e. spatially localised excitations that oscillate over time. These phenomena have subsequently been observed in various nonlinear optical waveguide arrays [1158804, b3727bba76384918a1c381ef76df660d], Bose-Einstein condensation in optical lattices [science.1062612, PhysRevLett.86.2353] and crystal lattice vibrations[PhysRevLett.82.3288]. In this type of system with two conserved quantities, namely energy and particle number, condensation into discrete breathers and a negative-temperature regime occur because the system exhibits an upper bound on energy [PhysRev.103.20, FLACH1998181], causing entropy to decrease as energy approaches this upper bound. The statistical mechanics analysis of the Discrete Nonlinear Schrödinger Equation (DNSE) for an isolated system was initially conducted by Rasmussen et al. [rasmussen_statistical_2000]. They showed that a transition from a disordered regime to a breather regime occurs at negative temperatures. Later, a statistical Gibbs measure was introduced to describe the localisation of breathers, by focusing on the matching condition between breather frequency and the chemical potential of the wave background, as proposed by Rumpf [rumpf_transition_2008, RuNe_PRL01, rumpf_growth_2007, rumpf_simple_2004, RUMPF20092067]. The dynamics of this system produces localisation of energy within a small number of isolated high-amplitude structures. In these publications, it was also argued that in the thermodynamic limit, states with several breathers will ultimately equilibrate to a single breather state, situated within a background of random fluctuations. However, numerical studies of nonequilibrium situations have shown that the relaxation process is extremely slow [PhysRevLett.134.097102, Ng_2009, PhysRevLett.122.084102, e19090445]. These authors give two main reasons for the slowness to equilibrium, one is that a system with localised states coarsens very slowly (entropic effect)[Iubini_2017, Iubini_2013], and the other is because of the dynamic decoupling of the high amplitude breathers from the background fluctuations (dynamical effects)[BaIuLiVu_PR21]. Recent papers [GrIuLiMa_JSM21, gradenigo2021condensation, giachello2025localization] report a phase transition in the condensation region of the negative-temperature regime of the DNSE, explaining the model’s thermodynamics in the microcanonical ensemble using large-deviation techniques, where the nearest-neighbour interaction is neglected, and more recently [giusfredi_mean-field_2025] by a mean-field theory that includes a simplified nearest-neighbour interaction.
@@
In this paper, we consider the model from the point of view of a canonical ensemble. Our formulation retains a parameter that measures the strength of the nonlinear term while the second constant of motion of the DNSE, the normalisation $N$, is taken to be fixed, $N=1$. Such a choice will turn out to be well-suited for our purpose, but one should keep in mind that a thermodynamic limit of large system size may differ from approaches where both the energy and the normalisation are considered as extensive variables. We set up a dissipative stochastic model, which follows by coupling the Hamiltonian of the DNSE to a microscopic heat bath (see appendix A for the details). Basic features of the resulting stochastic discrete nonlinear Schrödinger equation (SDNSE) are then discussed in section II, such as symmetries and general parameter dependencies (see also appendix B where the phase diagram of the related Hamiltonian system is discussed). In particular, we show that a change of sign in the coefficient of the nonlinear term effectively corresponds to a change of sign in temperature. As we have obtained our stochastic description from a consistent microscopic Hamiltonian model, the SDNSE obeys detailed balance and ensures relaxation towards a constrained canonical equilibrium distribution, which remains well defined even in the negative temperature regime (see appendix C for explicit calculations). Section III is devoted to a predominantly numerical study of the phase transition and its properties (some details of the numerical integration scheme are contained in appendix D). We begin with computing the dependence of the energy on the temperature, the so-called caloric equation of state. The dependence on the system parameters can be conveniently captured by a properly rescaled temperature variable. We find clear signatures of a first-order phase transition between a disordered and a localised phase. The transient nonequilibrium dynamics will display a non-monotonic dependence on the strength of the noise, a feature reminiscent of stochastic resonance. Finally, we provide in section IV an analytic approach for the phase transition which agrees quantitatively with the results of the simulations. We combine mean-field considerations and exact analytic results for the caloric equation of state of the linear model (see appendix E) to predict the temperature dependence of the energy in the disordered and in the localised phase, as well as the critical temperature itself. We also discuss how our findings relate to results available in the literature.
II The stochastic model
@@
We introduce here a class of stochastic models that describe the coupling of the Hamiltonian (1) to a heat bath with inverse temperature $1/(k_{B}T)=\beta$, where the normalization $N$ is still preserved (see e.g. [BoDe_CMP99, IuLeLiPo_JSM13, HaOl_SPDE23, IuPo_PRL25] for alternative stochastic models related to the DNSE). Such a class of models is given by the stochastic differential equations (see appendix A for a derivation of this model from first principles)
$\dot{c}_{k}(t)$ $=$ $i\frac{\partial H_{S}}{\partial\bar{c}_{k}}-\beta\frac{\sigma^{2}}{2}c_{k}(t)\left(\frac{\partial H_{S}}{\partial\bar{c}_{k}}\bar{c}_{k}(t)-\frac{\partial H_{S}}{\partial c_{k}}c_{k}(t)\right)$ (4)
$+i\sigma c_{k}(t)\xi_{k}(t)$
@@
II.1 Properties of the equations of motion
@@
$=$ $\{H_{S},N\}$
$-\beta\frac{\sigma^{2}}{2}\sum_{k=1}^{L}\left(|c_{k}|^{2}-|c_{k}|^{2}\right)\left(\frac{\partial H_{S}}{\partial\bar{c}_{k}}\bar{c}_{k}-\frac{\partial H_{S}}{\partial c_{k}}c_{k}\right)$
@@
$\dot{H}_{S}$ $=$ $\{H_{S},H_{S}\}$ (8)
$-\beta\frac{\sigma^{2}}{2}\sum_{k=1}^{L}\left(c_{k}\frac{\partial H_{S}}{\partial c_{k}}-\bar{c}_{k}\frac{\partial H_{S}}{\partial\bar{c}_{k}}\right)$
$\qquad\quad\quad\quad\left(\frac{\partial H_{S}}{\partial\bar{c}_{k}}\bar{c}_{k}-\frac{\partial H_{S}}{\partial c_{k}}c_{k}\right)$
@@
II.2 Symmetries and parameter dependencies
@@
The aforementioned symmetries can be demonstrated by looking at numerical simulations in the stationary regime. Figure 1 shows time traces of the energy $E(t)-2=H_{S}(t)-2$ and of $2-E(t)=2-H_{S}(t)$ for conjugate parameter settings, $(\alpha,\beta)$ and $(-\alpha,-\beta)$. The same symmetry can be seen when looking at the autocorrelation function of the energy.
Figure 1: (a) Time traces of the energy E(t)2E(t)-2 for parameter values $(\alpha,\beta)$ with positive $\alpha$ and of 2E(t)2-E(t) for conjugate parameter settings $(-\alpha,-\beta)$ obtained from numerical simulations of eq. (11) with system size $L=64$, noise strength $\sigma=0.3$ and stepsize $\tau=0.01$, generated from a localised initial condition $c_{n}(0)=\delta_{n,32}$ by skipping a transient of length about $\Delta t=1.6\times 10^{5}$ to ensure stationary behaviour. Conjugate parameter pairs: $(\alpha=16,\beta=10)$ (green), $(\alpha=-16,\beta=-10)$ (pink) and $(\alpha=-16,\beta=10)$ (blue), $(\alpha=16,\beta=-10)$ (orange). (b) The associated energy autocorrelation functions on a semi-logarithmic scale obtained from a time series of length about $T=2.6\times 10^{6}$, showing again the coincidence of data with conjugate parameter values. Parameter settings and colour coding as in part (a).
@@
Figure 2: (a) Time traces of the energy $E(t)$ for a system of size $L=64$ with parameter values $\alpha=16$, $\beta=10$ and different noise strength: $\sigma=0.01$ (green), $\sigma=0.03$ (orange), $\sigma=0.1$ (blue) and $\sigma=0.3$ (pink). Data have been obtained from a simulation of eq. (11) with stepsize $\tau=0.01$ and initial condition $c_{n}(0)=\delta_{n,32}$. The energy is shown as a function of the rescaled time variable $\sigma^{2}t$. The data collapse shows that $\sigma^{2}t$ sets the relevant time scale of the system. (b) Autocorrelation function of the energy in the stationary state, as a function of $\sigma^{2}t$ for the parameter setting (and colour coding) of part (a). Data have been computed from a time series of length about $T=2.6\times 10^{6}$ by skipping a transient of length about $\Delta t=16\times 10^{4}$. The deviations from exponential decay for small values of $\sigma$ are due to sampling errors in the numerical computation.
III Finite temperature phase transition
The Hamiltonian dynamics of the DNSE have been extensively discussed in the context of negative temperature states and related phase transition-like behaviour [rasmussen_statistical_2000, RuNe_PRL01, IuFrLiOpPo_NJP13, MiKaDaFl_PRL18]. Recently, a phase transition line has been reported [GrMa_JSM19, GrIuLiMa_JSM21] which shows a nontrivial system size dependence. This phase transition was found in the negative temperature regime based on analytical calculations in the microcanonical ensemble and stochastic simulations that strictly preserve energy and normalisation. With our choice of a negative sign of the nonlinearity in the Hamiltonian eq. (1), such a phase transition is expected to occur at positive temperatures. This allows us to uncover a phase transition in a setup when the DNSE is coupled to a heat bath using our stochastic model.
III.1 Caloric equation of state
@@
Figure 3: (a) Mean energy $E(\beta)$ computed from time traces of eq. (11) of length about $T=2.6\times 10^{6}$ (with longer time traces $T=10^{7}$ in the transition region) skipping a transient of length about $\Delta t=1.6\times 10^{5}$ for system size $L=64$, noise strength $\sigma=0.3$, and two different values of $\alpha$: $\alpha=8$ (green) and $\alpha=16$ (orange). Computations have been performed by an adiabatic parameter upsweep (open symbols) and downsweep (full symbols) of $\beta$, respectively. The black dashed line is the analytical result eq. (100) obtained for the case $\alpha=0$. (b) Variance of the energy for the same parameter setup (and colour coding) as in part (a).
@@
To obtain a better overview, we perform simulations for a larger range of parameter settings. Inspired by the partition integral, eqs. (1) and (10), which depends on the product $\alpha\beta/2$ when we discount for the interaction between sites, and by the particular system size dependence shown in eq. (100) it is tempting to introduce the rescaled temperature variable $\bar{\beta}=\alpha\beta/(2L)$. In fact, with such a rescaled variable, the phase transition seems to appear around $\bar{\beta}\approx 2.5$ for a larger set of parameters as indicated by a pronounced peak in the variance of the energy, see Figure 4(a). Below in section IV, we will support this observation with a simple mean-field model that yields a phase transition around $\bar{\beta}=2.45$. When increasing the system size, the phase transition is accompanied, as expected, by large time scales. That results in numerical hysteresis, reminiscent of first-order phase transitions, as shown in Figure 4(b). While increasing the noise strength seems to speed up the dynamics, as suggested by the results of the previous section, the time scales become too long to be covered by standard numerical simulation times, and only quasi-stationary behaviour can be realised.
@@
Figure 4: (a) Variance of the energy in dependence on the rescaled temperature variable $\bar{\beta}=\alpha\beta/(2L)$ for $\sigma=0.3$ and $L=32$ (open symbols), $L=64$ (full symbols), $\alpha=8$ (green), $\alpha=16$ (orange), $\alpha=32$ (blue), showing that the location of the phase transition depends essentially on the parameter $\bar{\beta}=\beta\alpha/2L$. Data have been computed by adiabatic parameter upsweeps (solid) and downsweeps (dashed) taking time averages over an interval of length about $T=2.6\times 10^{6}$ while skipping a transient of length about $\Delta t=1.6\times 10^{5}$. The three black points indicate the particular temperature values used in Figure 5. (b) Mean energy $E(\bar{\beta})$ for a system of size $L=128$ (green) and $L=64$ (orange) with $\alpha=16$ and $\sigma=0.3$, showing hysteresis due to extensive equilibration times for larger system sizes. Data have been computed by adiabatic parameter upsweeps (full symbols, solid) and downsweeps (open symbols, dashed) taking time averages over an interval of length about $T=2.6\times 10^{6}$ while skipping a transient of length about $\Delta t=1.6\times 10^{4}$. The black dash-dotted line is the equilibrium energy of the system predicted by equation (12), with the blue dashed line indicating analytic estimates of metastable states (see section IV for details).
III.2 Localised and disordered phase

### v2

@@ Abstract
keywords:
DNSE, pattern formation, soliton, phase transition
articletype: Paper
[…]
1 Introduction
@@
The one-dimensional discrete nonlinear Schrödinger equation (DNSE) is a Hamiltonian system with two conserved quantities that exhibits intriguing features such as localised modes (breathers) and negative temperature states [BaIuLiVu_PR21]. This non-integrable model was initially proposed to describe the dynamics of a group of anharmonic oscillators. Its primary aim was to explain nonlinear localisation phenomena, specifically breathers, i.e. spatially localised excitations that oscillate over time. These phenomena have subsequently been observed in various nonlinear optical waveguide arrays [1158804, b3727bba76384918a1c381ef76df660d], Bose-Einstein condensation in optical lattices [science.1062612, PhysRevLett.86.2353] and crystal lattice vibrations[PhysRevLett.82.3288]. In this type of system with two conserved quantities, namely energy and particle number, condensation into discrete breathers and a negative-temperature regime occur because the system exhibits an upper bound on energy [PhysRev.103.20, FLACH1998181], causing entropy to decrease as energy approaches this upper bound. The statistical mechanics analysis of the Discrete Nonlinear Schrödinger Equation (DNSE) for an isolated system was initially conducted by Rasmussen et al. [rasmussen_statistical_2000]. They showed that a transition from a disordered regime to a breather regime occurs at negative temperatures. Later, Rumpf, through a series of publications, identified the probability distribution of breathers and demonstrated that the entropy as a function of energy exhibits nonanalytic behaviour at the transition [rumpf_transition_2008, RUMPF20092067]. Furthermore, Rumpf established that the focusing process, characterised by the emergence and amplification of peaks, is driven by entropy generation in the state of small amplitude waves resulting from the two conserved quantities of the system [rumpf_growth_2007, rumpf_simple_2004, RuNe_PRL01]. The dynamics of this system produces localisation of energy within a small number of isolated high-amplitude structures. In these publications, it was also argued that in the thermodynamic limit, states with several breathers will ultimately equilibrate to a single breather state, situated within a background of random fluctuations. However, numerical studies of nonequilibrium situations have shown that the relaxation process is extremely slow [PhysRevLett.134.097102, Ng_2009]. These authors give two main reasons for the slowness to equilibrium [PhysRevLett.122.084102, e19090445], one is that a system with localised states coarsens very slowly (entropic effect)[Iubini_2017, Iubini_2013], and the other is because of the dynamic decoupling of the high amplitude breathers from the background fluctuations (dynamical effects)[BaIuLiVu_PR21]. Recent papers [GrIuLiMa_JSM21, gradenigo2021condensation, giachello2025localization] report a phase transition in the condensation region of the negative-temperature regime of the DNSE, explaining the model’s thermodynamics in the microcanonical ensemble using large-deviation techniques, where the nearest-neighbour interaction is neglected, and more recently [giusfredi_mean-field_2025] by a mean-field theory that includes a simplified nearest-neighbour interaction.
@@
In this paper, we consider the model from the point of view of a canonical ensemble. Our formulation retains a parameter that measures the strength of the nonlinear term while the second constant of motion of the DNSE, the normalisation $N$, is taken to be fixed, $N=1$. Such a choice will turn out to be well-suited for our purpose, but one should keep in mind that a thermodynamic limit of large system size may differ from approaches where both the energy and the normalisation are considered as extensive variables. We set up a dissipative stochastic model, which follows by coupling the Hamiltonian of the DNSE to a microscopic heat bath (see appendix A for the details). Basic features of the resulting stochastic discrete nonlinear Schrödinger equation (SDNSE) are then discussed in section 2, such as symmetries and general parameter dependencies (see also appendix B where the phase diagram of the related Hamiltonian system is discussed). In particular, we show that a change of sign in the coefficient of the nonlinear term effectively corresponds to a change of sign in temperature. As we have obtained our stochastic description from a consistent microscopic Hamiltonian model, the SDNSE obeys detailed balance and ensures relaxation towards a constrained canonical equilibrium distribution, which remains well defined even in the negative temperature regime (see appendix C for explicit calculations). Section 3 is devoted to a predominantly numerical study of the phase transition and its properties (some details of the numerical integration scheme are contained in appendix D). We begin with computing the dependence of the energy on the temperature, the so-called caloric equation of state. The dependence on the system parameters can be conveniently captured by a properly rescaled temperature variable. We find clear signatures of a first-order phase transition between a disordered and a localised phase. The transient nonequilibrium dynamics will display a non-monotonic dependence on the strength of the noise, a feature reminiscent of stochastic resonance. Finally, we provide in section 4 an analytic approach for the phase transition which agrees quantitatively with the results of the simulations. We combine mean-field considerations and exact analytic results for the caloric equation of state of the linear model (see appendix E) to predict the temperature dependence of the energy in the disordered and in the localised phase, as well as the critical temperature itself. We also discuss how our findings relate to results available in the literature.
2 The stochastic model
@@
We introduce here a class of stochastic models that describe the coupling of the Hamiltonian (1) to a heat bath with inverse temperature $1/(k_{B}T)=\beta$, where the normalization $N$ is still preserved (see e.g. [BoDe_CMP99], for early alternative stochastic models and [IuLeLiPo_JSM13, HaOl_SPDE23, IuPo_PRL25] for more recent developments of the stochastic models related to the DNSE). Such a class of models is given by the stochastic differential equations (see appendix A for a derivation of this model from first principles)
$\dot{c}_{k}(t)=i\frac{\partial H_{S}}{\partial\bar{c}_{k}}-\beta\frac{\sigma^{2}}{2}c_{k}(t)\left(\frac{\partial H_{S}}{\partial\bar{c}_{k}}\bar{c}_{k}(t)-\frac{\partial H_{S}}{\partial c_{k}}c_{k}(t)\right)+i\sigma c_{k}(t)\xi_{k}(t)$ (4)
@@
2.1 Properties of the equations of motion
@@
$=$ $\{H_{S},N\}-\beta\frac{\sigma^{2}}{2}\sum_{k=1}^{L}\left(|c_{k}|^{2}-|c_{k}|^{2}\right)\left(\frac{\partial H_{S}}{\partial\bar{c}_{k}}\bar{c}_{k}-\frac{\partial H_{S}}{\partial c_{k}}c_{k}\right)$
@@
$\dot{H}_{S}$ $=$ $\{H_{S},H_{S}\}-\beta\frac{\sigma^{2}}{2}\sum_{k=1}^{L}\left(c_{k}\frac{\partial H_{S}}{\partial c_{k}}-\bar{c}_{k}\frac{\partial H_{S}}{\partial\bar{c}_{k}}\right)\left(\frac{\partial H_{S}}{\partial\bar{c}_{k}}\bar{c}_{k}-\frac{\partial H_{S}}{\partial c_{k}}c_{k}\right)$ (8)
@@
2.2 Symmetries and parameter dependencies
@@
The aforementioned symmetries can be demonstrated by looking at numerical simulations in the stationary regime. Figure 1 shows time traces of the energy $E(t)-2=H_{S}(t)-2$ and of $2-E(t)=2-H_{S}(t)$ for conjugate parameter settings, $(\alpha,\beta)$ and $(-\alpha,-\beta)$. The same symmetry can be seen when looking at the autocorrelation function of the energy, $C_{EE}(t)=\langle E(t)E(0)\rangle-\langle E(0)\rangle^{2}$, where the average $\langle.\rangle$ is taken over a stationary ensemble of initial conditions $E(0)$, which can for instance be realised by a long time average. For practical purposes the autocorrelation function is computed via the Fourier transfrom of a stationary time series, using the Wiener Khinchin theorem.
Refer to captionRefer to caption
Figure 1: (a) Time traces of the energy E(t)2E(t)-2 for parameter values $(\alpha,\beta)$ with positive $\alpha$ and of 2E(t)2-E(t) for conjugate parameter settings $(-\alpha,-\beta)$ obtained from numerical simulations of eq. (11) with system size $L=64$, noise strength $\sigma=0.3$ and stepsize $\tau=0.01$, generated from a localised initial condition $c_{n}(0)=\delta_{n,32}$ by skipping a transient of length about $\Delta t=1.6\times 10^{5}$ to ensure stationary behaviour. Conjugate parameter pairs: $(\alpha=16,\beta=10)$ (green), $(\alpha=-16,\beta=-10)$ (pink) and $(\alpha=-16,\beta=10)$ (blue), $(\alpha=16,\beta=-10)$ (orange). (b) The associated energy autocorrelation functions on a semi-logarithmic scale obtained from a time series of length about $T_{sim}=2.6\times 10^{6}$, showing again the coincidence of data with conjugate parameter values. Parameter settings and colour coding as in part (a).
@@
Refer to captionRefer to caption
Figure 2: (a) Time traces of the energy $E(t)$ for a system of size $L=64$ with parameter values $\alpha=16$, $\beta=10$ and different noise strength: $\sigma=0.01$ (green), $\sigma=0.03$ (orange), $\sigma=0.1$ (blue) and $\sigma=0.3$ (pink). Data have been obtained from a simulation of eq. (11) with stepsize $\tau=0.01$ and initial condition $c_{n}(0)=\delta_{n,32}$. The energy is shown as a function of the rescaled time variable $\sigma^{2}t$. The data collapse shows that $\sigma^{2}t$ sets the relevant time scale of the system. (b) Autocorrelation function of the energy in the stationary state, as a function of $\sigma^{2}t$ for the parameter setting (and colour coding) of part (a). Data have been computed from a time series of length about $T_{sim}=2.6\times 10^{6}$ by skipping a transient of length about $\Delta t=16\times 10^{4}$. The deviations from exponential decay for small values of $\sigma$ are due to sampling errors in the numerical computation.
3 Finite temperature phase transition
The Hamiltonian dynamics of the DNSE have been extensively discussed in the context of negative temperature states [IuFrLiOpPo_NJP13] and related phase transition-like behaviour [rasmussen_statistical_2000, RuNe_PRL01, MiKaDaFl_PRL18]. Recently, a phase transition line has been reported [GrMa_JSM19, GrIuLiMa_JSM21] which shows a nontrivial system size dependence. This phase transition was found in the negative temperature regime based on analytical calculations in the microcanonical ensemble and stochastic simulations that strictly preserve energy and normalisation. With our choice of a negative sign of the nonlinearity in the Hamiltonian eq. (1), such a phase transition is expected to occur at positive temperatures. This allows us to uncover a phase transition in a setup when the DNSE is coupled to a heat bath using our stochastic model.
3.1 Caloric equation of state
@@
Figure 3: (a) Mean energy $E(\beta)$ computed from time traces of eq. (11) of length about $T_{sim}=2.6\times 10^{6}$ (with longer time traces $T_{sim}=10^{7}$ in the transition region) skipping a transient of length about $\Delta t=1.6\times 10^{5}$ for system size $L=64$, noise strength $\sigma=0.3$, and two different values of $\alpha$: $\alpha=8$ (green) and $\alpha=16$ (blue). Computations were performed using an adiabatic parameter upsweep (open symbols) and downsweep (full symbols) of $\beta$; the data for upsweep and downsweep coincide perfectly. The black dashed line is the analytical result eq. (99) obtained for the case $\alpha=0$. (b) Variance of the energy for the same parameter setup (and colour coding) as in part (a).
@@
To obtain a better overview, we perform simulations for a larger range of parameter settings. Inspired by the partition integral, eqs. (1) and (10), which depends on the product $\alpha\beta/2$ when we discount for the interaction between sites, and by the particular system size dependence shown in eq. (99) it is tempting to introduce the rescaled inverse temperature variable $\bar{\beta}=\alpha\beta/(2L)$. In fact, with such a rescaled variable, the phase transition seems to appear around $\bar{\beta}\approx 2.5$ for a larger set of parameters as indicated by a pronounced peak in the variance of the energy, see Figure 4(a). Below in section 4, we will support this observation with a simple mean-field model that yields a phase transition around $\bar{\beta}=2.45$. When increasing the system size, the phase transition is accompanied, as expected, by large time scales. That results in numerical hysteresis, reminiscent of first-order phase transitions, as shown in Figure 4(b). Increasing the noise strength seems to accelerate the system dynamics as suggested by the results of the previous section (for a more detailed discussion of the impact of noise we refer the reader to section 3.2)). Therefore increasing the noise strength partially suppresses numerical hysteresis, see Figure 4(b). Having said that, the time scales may become too long to be covered by standard numerical simulation times, and only quasi-stationary behaviour can be realised.
Refer to captionRefer to caption
@@
Figure 4: (a) Variance of the energy in dependence on the rescaled inverse temperature variable $\bar{\beta}=\alpha\beta/(2L)$ for $\sigma=0.3$ and $L=32$ (triangle symbols/lighter colours), $L=64$ (square symbols/darker colours), $\alpha=8$ (green), $\alpha=16$ (blue), $\alpha=32$ (red), showing that the location of the phase transition depends essentially on the parameter $\bar{\beta}=\beta\alpha/2L$. Data have been computed by adiabatic parameter upsweeps (solid) and downsweeps (dashed) taking time averages over an interval of length about $T_{sim}=2.6\times 10^{6}$ while skipping a transient of length about $\Delta t=1.6\times 10^{5}$. The three black points indicate the particular temperature values used in Figure 5. (b) Mean energy $E(\bar{\beta})$ for a system of size $L=128$ (dark blue circles) and $L=64$ (blue squares) with $\alpha=16$ and $\sigma=0.3$, showing hysteresis due to extensive equilibration times for larger system sizes. Data have been computed by adiabatic parameter upsweeps (open symbols, solid) and downsweeps (full symbols, dashed) taking time averages over an interval of length about $T_{sim}=2.6\times 10^{6}$ while skipping a transient of length about $\Delta t=1.6\times 10^{4}$. The red dash line is the equilibrium energy of the system predicted by equation (12), with the orange dashed-dotted line indicating analytic estimates of metastable states (see section 4 for details).
3.2 Localised and disordered phase